OpenStudy (anonymous):

A Web music store offers two versions of a popular song. The size of the standard version is 2.7 megabytes (MB). The size of the high-quality version is 4.5 MB. Yesterday, the high-quality version was downloaded twice as often as the standard version. The total size downloaded for the two versions was 2574 MB. How many downloads of the standard version were there?

OpenStudy (chosenmatt):

need help? :)

OpenStudy (anonymous):

2.7y+4.5 x=2574 MB

OpenStudy (anonymous):

then solve for y and x?

OpenStudy (anonymous):

you need other eq

OpenStudy (anonymous):

x hight q y low q x=2y

OpenStudy (anonymous):

now you can slove

OpenStudy (anonymous):

don't understand the x hight q y low q

OpenStudy (anonymous):

2.7y+4.5 x=2574 MB 2.7y+4.5(2y)=2574 MB

OpenStudy (anonymous):

y number of standard version x number of high q

OpenStudy (anonymous):

solve for y